Subtract 4p2q – 3pq + 5pq2 – 8p + 7q – 10 from 18 – 3p – 11q + 5pq – 2pq^2 + 5p^2q

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(18 - 3p - 11q + 5pq - 2pq^2+ 5p^2q) - (4p^2q - 3pq + 5pq^2- 8p + 7q - 10)
= 18 - 3p - 11q + 5pq - 2pq^2+ 5p^2q - 4p^2q + 3pq - 5pq^2+ 8p - 7q + 10
=18+10 -3p+8p -11q – 7q + 5 pq+ 3pq- 2pq^2 – 5pq^2 + 5 p^2 q – 4p^2 q
= 28 + 5p - 18q + 8pq - 7pq^2+ p^2q

Thanks